OpenStudy (anonymous):

1. Riley noticed an aquarium in his father’s office. The aquarium had the dimensions 16 in. by 8.5 in. by 10.5 in. The formula for volume is: V = l × w × h. (a) Riley noticed that the aquarium was three-fourths full of water. How many more cubic inches of water would be required to fill the tank? Show your work. (b) Another aquarium in the building has dimensions that are each triple the size of the aquarium in Riley’s father’s office. Riley thought that the volume would also triple. Is Riley correct? How many times greater is the volume of the larger aquarium than the smaller?

OpenStudy (danjs):

First off, find the volume of the tank [in^3] by multiplying together the 3 given numbers

OpenStudy (danjs):

Volume = length x width x height length = 16 width = 8.5 height = 10.5

OpenStudy (anonymous):

1428

OpenStudy (danjs):

now find 3/4ths of that, multiply it by 0.75

OpenStudy (anonymous):

1428/4=357

OpenStudy (danjs):

close... \[1428 *\frac{ 3 }{ 4 } = \frac{ 1428 * 3 }{ 4 }\]

OpenStudy (anonymous):

1071

OpenStudy (danjs):

Right, then 3/4ths of the tank is 1071 in^3 of water. How much is missing?

OpenStudy (danjs):

Full volume 1428 - 3/4ths volume 1071 = 357 in^3 left to fill the tank

OpenStudy (danjs):

Or you could have said, 1/4th of the water is missing, and did what you did above, and said 1/4th of 1428 is 357 in^3 of water

OpenStudy (danjs):

Do you know how to begin the second part?

OpenStudy (anonymous):

16*3

OpenStudy (danjs):

right, multiply all 3 of those dimensions by 3

OpenStudy (danjs):

Larger Tank Length = 16*3 = 48 in Width = 8.5*3 = 25.5 in height = 10.5*3 = 31.5 in

OpenStudy (anonymous):

38,556

OpenStudy (danjs):

How many smaller tanks will fit in the larger tank?

OpenStudy (anonymous):

38,556-4284=34272

OpenStudy (danjs):

you don't want to subtract the two values, you want to see how many times 4284 will go into 38556, divide the larger by the smaller

OpenStudy (anonymous):

ooo

OpenStudy (anonymous):

9

OpenStudy (danjs):

or you could keep subtracting , and count how many times it takes subtracting to get to zero

OpenStudy (danjs):

Here is a summary of what you did...

OpenStudy (danjs):

wait

OpenStudy (anonymous):

ok

OpenStudy (danjs):

The smaller volume v = 16 * 8.5 * 10.5 = 1428 The Larger Volume V = 48 * 25.5 * 31.5 = 38556 The Larger Volume divided by the smaller volume 38556 / 1428 = 27 The larger tank is 27 times the volume of the smaller tank

OpenStudy (danjs):

What you basically did is... Small volume : v = l x w x h Large Volume V = 3l x 3w x 3h = (3x3x3)*(l x w x h) = 27

OpenStudy (danjs):

If you multiply the side lengths by 3, then the volume will be 3^3 or 27 times larger

OpenStudy (danjs):

In general If you multiply the lengths by a factor of 'n' the Larger volume will be n^3 times larger than the smaller volume

OpenStudy (anonymous):

ok

OpenStudy (danjs):

one of the numbers was wrong, you used the value from the first part, the 1/4th of the tank 1071 in the second part, we needed to use the whole volume 1428

OpenStudy (danjs):

sorry it is late, sleepy and missed it

OpenStudy (danjs):

feel free to mssg me and fan or whatever, i am usually on once a day at at least... ttyl

OpenStudy (anonymous):

ok thanks for the help

OpenStudy (danjs):

welcome